40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Thu Jun 30 16:11:57 2016, ericq, Update, General, Vent Prep 
    Reply  Thu Jun 30 21:24:36 2016, ericq, Update, General, Vent Prep 
       Reply  Fri Jul 1 09:20:44 2016, Steve, Update, VAC, vent 78 vent78.pngvent78at250T.png
          Reply  Fri Jul 1 15:48:10 2016, Steve, Update, VAC, vent 78 at 710 Torr pd78at710T.png
             Reply  Wed Jul 6 10:58:07 2016, Steve, Update, VAC, vent 78 valve configuration CamberOpen.png
                Reply  Wed Jul 20 11:39:07 2016, Steve, Update, VAC, vent 78 day 20 vent78-d20.png
                   Reply  Mon Aug 8 10:35:12 2016, Steve, Update, PEM, vent 78 day 38 Vent38days.pngvent78d40RgaBg.png
                      Reply  Tue Oct 4 11:02:31 2016, Steve, Update, VAC, RGA is out of order lastBGscan.png
                         Reply  Fri Oct 7 10:29:57 2016, Steve, Update, VAC, RGA removed RGAremoved.png
                            Reply  Mon Oct 31 16:20:50 2016, Steve, Update, VAC, RGA is reinstalled 
                            Reply  Tue Nov 22 11:31:52 2016, Steve, Update, VAC, RGA is running again RGAisBack33d.pngpd80-560Hz-d33.pngafterRepair1d.png
                               Reply  Mon Nov 28 09:45:44 2016, Steve, Update, VAC, RGA scan at pumpdown day 40 RGAscan_d40.pngd2_&_d7scan.pngRGAd2.png
                                  Reply  Thu Dec 22 09:37:20 2016, Steve, Update, VAC, RGA scan at day 63 pd80-d63.pngpd80-580Hz-d63.png
                                     Reply  Tue Jan 10 14:24:09 2017, Steve, Update, VAC, RGA scan at day 82 pd80VNd82.png
                                        Reply  Mon Feb 27 15:33:10 2017, Steve, Update, VAC, RGA scan at day 130 rgascan@130d.png
          Reply  Wed Jul 6 08:34:00 2016, Steve, Update, VAC, RGA scan at day 639 RGAscan639d.pngpd78d640.png
          Reply  Mon Sep 12 17:54:24 2016, Steve, Update, VAC, pumpdown 79 
Message ID: 12483     Entry time: Mon Sep 12 17:54:24 2016     In reply to: 12237
Author: Steve 
Type: Update 
Category: VAC 
Subject: pumpdown 79 

 We are pumping down. The annuloses are below 10 mTorr

The vacuum gauges are not communicating with the medm so there is no plot available.

The main volume pumping is stopped at P1 = 220 Torr  ( New SuperBee 174 Torr ) for overnight.

note: SuperBee was reading 791 Torr at atm and it was not set to 760 !

ELOG V3.1.3-